OpenStudy (anonymous):

Please help, What is the value for X in the proportion X+8/5x-2 = 3/8

OpenStudy (anonymous):

What @stamp

OpenStudy (anonymous):

(x + 8) / (5x - 2) = 3/8 Is this the problem ?

OpenStudy (anonymous):

Yes @kelliegirl33

OpenStudy (anonymous):

So what's the value for x then ?

OpenStudy (anonymous):

(x + 8)/(5x - 2) = 3/8 cross multiply 8(x + 8) = 3(5x - 2) 8x + 64 = 15x - 6 8x - 15x = -6 -64 -7x = -70 x = 10 check.. (x + 8)/(5x - 2) = 3/8 (10 + 8) / (5(10) - 2) = 3/8 18/48 = 3/8 3/8 = 3/8 (correct)

OpenStudy (anonymous):

Thank u @kelliegirl33 Which of the following represents The linear equation 2x=y+3 in slope intercept form? 2X + Y = 3 2X - Y = 3 Y=2X-3 Y=2x+3

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

2x = y + 3 2x - 3 = y or y = 2x - 3

OpenStudy (anonymous):

slope intercept is y = mx + b....so your answer is y = 2x - 3